One canned juice drink is 15% orange juice, another is 5% orange juice. How many liters of each should be mixed together in order to get 10L that is 6% orange juice?
How many liters of the 15% orange juice should be in the mixture?

Answer by mananth(16946)

Percent ---------------- quantity
Juice I 15.00% ---------------- x liters
Juice II 5.00% ------ 10 - x liters
Mixture 6.00% ---------------- 10
Total 10 liters
15.00% x + 5.00% ( 10 - x ) = 6.00% * 10
15 x + 5 ( 10 - x ) = 60
15 x + 50 - 5 x = 60
15 x - 5 x = 60 - -50
10 x = 10
/ 10
x = 1 liter 15.00% Juice I
9 liters 5.00% Juice II

m.ananth@hotmail.ca